### v3

# MM-SAP: A Comprehensive Benchmark for Assessing Self-Awareness of Multimodal Large Language Models in Perception

Abstract
Recent advancements in Multimodal Large Language Models (MLLMs) have demonstrated exceptional capabilities in visual perception and understanding. However, these models also suffer from hallucinations, which limit their reliability as AI systems. We believe that these hallucinations are partially due to the models’ struggle with understanding what they can and cannot perceive from images, a capability we refer to as self-awareness in perception. Despite its importance, this aspect of MLLMs has been overlooked in prior studies. In this paper, we aim to define and evaluate the self-awareness of MLLMs in perception. To do this, we first introduce the knowledge quadrant in perception, which helps define what MLLMs know and do not know about images. Using this framework, we propose a novel benchmark, the Self-Awareness in Perception for MLLMs (MM-SAP), specifically designed to assess this capability. We apply MM-SAP to a variety of popular MLLMs, offering a comprehensive analysis of their self-awareness and providing detailed insights. The experiment results reveal that current MLLMs possess limited self-awareness capabilities, pointing to a crucial area for future advancement in the development of trustworthy MLLMs. Code and data are available at https://github.com/YHWmz/MM-SAP.
MM-SAP: A Comprehensive Benchmark for Assessing Self-Awareness of Multimodal Large Language Models in Perception
Yuhao Wang1, Yusheng Liao1, Heyang Liu1, Hongcheng Liu1, Yanfeng Wang1,2, Yu Wang1,2🖂 1Cooperative Medianet Innovation Center, Shanghai Jiao Tong University 2Shanghai Artificial Intelligence Laboratory {colane,liao20160907,liuheyang,hongcheng_liu,wangyanfeng622,yuwangsjtu}@sjtu.edu.cn
footnotemark: footnotetext: 🖂:Corresponding author.
1 Introduction
Recently, breakthrough advances in large language models (LLMs) have greatly reshaped the artificial intelligence landscape  Brown et al. (2020); Chowdhery et al. (2023); Touvron et al. (2023); OpenAI (2023a); Bubeck et al. (2023). Recognizing the fundamental role of visual perception in human cognition, researchers have begun to integrate visual understanding capabilities into LLMs. This integration has led to the emergence of Multimodal Large Language Models (MLLMs) Yin et al. (2023a); Zhang et al. (2024). Early works expanded the capabilities by incorporating visual encoders Zhu et al. (2023); Dai et al. (2023); Liu et al. (2023c), thus enabling them to recognize image content. Subsequent developments, exemplified by GPT-4V OpenAI (2023b) and Gemini Team et al. (2023), have further demonstrated the immense potential of MLLMs.
Figure 1: Self-awareness of a trustworthy MLLM. A trustful MLLM can be aware of what it knows and what it does not know. Top: For the questions it knows, it would provide correct answers as a reliable AI system. Bottom: It can recognize unknown questions and refuse to give answers, preventing the generation of incorrect responses.
Despite their impressive vision-language understanding capabilities, MLLMs are not yet considered trustworthy AI systems Li et al. (2023a). Prior researches have shown that these models can generate inconsistent responses to input images, a phenomenon often referred to as ‘hallucination’ Liu et al. (2023a); Li et al. (2023c). A key reason for this is the MLLMs’ limited self-awareness, meaning their understanding of what they know and what they do not know. This gap in self-awareness often leads to overconfidence in their outputs, regardless of whether the generated content matches the images or not. Enhancing MLLMs’ ability to recognize their own limitations is essential for enabling them to accurately determine when to express uncertainty and limitation in their responses, thereby avoiding hallucinations. Previous studies have investigated the self-awareness of LLMs Yin et al. (2023b); Amayuelas et al. (2023). These studies categorize the knowledge of LLMs using the knowledge quadrant shown in Figure 2(a), and explore how LLMs respond to unknown questions.  Cheng et al. (2024) further constructed an ‘Idk’ dataset to enhance LLMs’ self-awareness, resulting in more truthful AI assistants. However, these studies have not explored the self-awareness of MLLMs, which is more complex than that of LLMs due to the multimodal inputs.
In this paper, we delve into the pivotal role of self-awareness in image perception for MLLMs, underscoring its importance for the creation of trustworthy AI systems. Self-awareness, the ability of MLLMs to assess their own information boundaries, enabling them to deliver reliable responses while acknowledging their limitations. This capability ensures that MLLMs can provide precise answers when confident and, crucially, refrain from offering responses when the query surpasses their understanding or the visual information provided (Figure 1). Recognizing the insufficiency of existing frameworks, which are primarily tailored to unimodal LLMs, our work first introduces an expanded knowledge quadrant that incorporates visual inputs, offering a more nuanced and comprehensive approach to defining and evaluating self-awareness for MLLMs in image perception. This innovative quadrant, illustrated in Figure 2(b), is specifically designed to address the complexities and challenges inherent in multimodal scenarios. By systematically mapping out the landscape of knowns and unknowns in the context of visual perception, our proposed knowledge quadrant lays the foundation for assessing and enhancing the reliability and trustworthiness of MLLMs.
Furthermore, leveraging the proposed Knowledge Quadrant for MLLMs, we design and introduce the Self-Awareness in Perception for MLLMs (MM-SAP) benchmark, a dataset designed to specifically evaluate MLLMs’ self-awareness in perception. MM-SAP stands out by assessing both the models’ ability to interpret visual information and the recognition of their limitations, marking a significant difference from existing benchmarks. This dual-focus evaluation provides a holistic view of MLLMs’ self-awareness capabilities. Our extensive evaluation of thirteen prominent MLLMs using MM-SAP has yielded insightful findings, showcasing how these models manage their knowledge boundaries.In summary, our main contributions are as follows:
(a) Knowledge Quadrant for LLMs
(b) Knowledge Quadrant for MLLMs
Figure 2: Knowledge quadrants for LLMs and MLLMs. Taking the visual information into account, we expand the original quadrant horizontally to develop the knowledge quadrant for MLLMs.
  • Developing the Knowledge Quadrant for MLLMs: We propose a novel framework, the Knowledge Quadrant for MLLMs, designed to enhance our understanding of self-awareness in MLLMs. This framework innovatively incorporates visual perception into the assessment of MLLMs’ self-awareness, offering a structured approach to examining how these models process and interpret multimodal information. It lays the groundwork for future advancements in improving self-awareness in MLLMs and creating more trustworthy MLLMs.
  • A Pioneering Benchmark for MLLM Evaluation: The MM-SAP dataset we introduce in this paper serves as a novel benchmark for evaluating the self-awareness of MLLMs, specifically in their ability to perceive and interpret visual information. This benchmark is designed to test MLLMs on their recognition of what they know and what they do not know, providing a crucial tool for this field. MM-SAP stands out for its focus on both knowns and unknowns, facilitating a deeper understanding of where MLLMs excel and where they fall short, thereby guiding future enhancements in model development.
  • Comprehensive Assessment of MLLMs’ Self-Awareness Capabilities: Our evaluation of thirteen prominent MLLMs using the MM-SAP benchmark yields insightful results regarding the current capabilities of MLLMs in terms of self-awareness. While these models show competence in dealing with information within their knowledge base, they often falter in recognizing the limits of their understanding. This analysis highlights a vital area for improvement in MLLM research, suggesting a clear need for strategies that bolster models’ ability to identify and acknowledge their informational boundaries.
2 Related work
2.1 Self-awareness of LLMs
Previous works have explored LLMs’ self-awareness, assessing their abilities to recognize their limitations.  Amayuelas et al. (2023) collected a dataset named the Known-Unknown Questions (KUQ) to assess the LLMs’ ability to classify known and unknown questions.  Yin et al. (2023b) introduced SelfAware, comprising unanswerable questions and their answerable counterparts, to evaluate the uncertainty in LLM’s responses.  Cheng et al. (2024) aligned AI assistants with an ’I don’t know’ (Idk) dataset which contains both known and unknown questions, enhancing their reliability. Distinct from these endeavors, our work pioneers the exploration of self-awareness within the context of multimodal scenarios, addressing a critical gap in existing research.
2.2 Hallucination on MLLMs
For MLLMs, hallucinations are generally defined as situations where the generated responses contain information that is not present in the image Cui et al. (2023). Previous studies have purposed various dataset to assess the hallucinations of MLLMs Wang et al. (2023a); Cui et al. (2023); Li et al. (2023b); Guan et al. (2023). To alleviate this problem,  Liu et al. (2023a) developed a balanced instructions datasets comprising both positive and negative samples.  Yu et al. (2023a) proposed RLHF-V to enhances MLLM trustworthiness. However, the connection between MLLMs’ self-awareness and hallucinations remains unexplored. Our work addresses this gap by proposing the Knowledge Quadrant for MLLMs and the MM-SAP, marking a novel direction in improving self-awareness to mitigate hallucination.
2.3 Benchmarks for MLLMs
The evolution of MLLMs has spurred the development of benchmarks like MME Fu et al. (2023), MMBench Liu et al. (2023d), MM-Vet Yu et al. (2023b), and MathVista Lu et al. (2023), each designed to assess various aspects of MLLM performance. These benchmarks have significantly advanced our understanding of MLLMs’ perceptual, cognitive, and reasoning capabilities. Distinctively, our works introduce a novel focus on evaluating MLLMs’ self-awareness, emphasizing the critical need for MLLMs to recognize what they know and what they do not. This marks a pivotal step towards developing more reliable and trustworthy MLLMs.
3 Self-awareness in Perception
Self-awareness refers to a model’s ability to recognize its information limitations, encompassing their capabilities to discern ‘knowns’ and ‘unknowns’. For LLMs, we can categorize their knowledge using the knowledge quadrant framework to evaluate their self-awareness. However, this framework encounters great complexity when applied to MLLMs due to the inclusion of visual inputs. In this work, we narrow our focus to self-awareness in image perception, namely, the ability of MLLMs to recognize the information they can and cannot perceive from images.
Figure 3: Overview of MM-SAP. Our MM-SAP benchmark comprises three sub-datasets, namely BasicVisQA, KnowVisQA, and BeyondVisQA, and includes a total of 19 subtasks. The white dashed line indicates that the delineation between ‘Knowns’ and ‘Unknowns’ is model-specific. The number in square brackets in the middle ring represents the size of the subset, while the number in the outer ring indicates the proportion of each subtask within the subset.
3.1 Knowledge Quadrant for MLLMs
We first divide perceptual questions into two categories: those answerable based on image information and those querying information not present in the image (e.g., non-existent objects). The latter is always beyond the reach of MLLMs as they cannot access the necessary information. We further classify the answerable questions on the need for knowledge to provide an answer. For perceptual questions that do not require external knowledge, such as those concerning object attributes, MLLMs need to extract basic visual information like color or shape from images. We suggest that MLLMs have learned these basic visual concepts through multimodal instruction tuning. Consequently, we believe MLLMs possess sufficient information to address these questions. However, there are instances where MLLMs need visual knowledge to recognize image content, such as brand and landmark recognition. Whether MLLMs can answer these questions depends on the models’ knowledge boundaries.
Therefore, to develop the knowledge quadrant for MLLMs, we need to consider not only the intrinsic knowledge within model parameters, but also the external information provided by images in multimodal scenario. Based on the above analysis, we categorize information in image perception into three types: basic visual information, knowledge-intensive visual information, and information beyond the input images. We classify both basic visual information and the model’s inherent visual knowledge as ‘knowns’, whereas visual information that lies beyond the image and the model’s unknown visual knowledge is categorized as ‘unknowns’. In light of this categorization, we consider visual information in our analysis, describe ‘knowns’ and ‘unknowns’ for MLLMs in the context of image perception, and further introduce a knowledge quadrant specifically tailored for MLLMs, as shown in Figure 2(b).
The knowledge quadrant categorizes information in image perception into four segments: Known Knowns, Known Unknowns, Unknown Knowns, and Unknown Unknowns.Known Knowns are information that models know and are aware of knowing. In contrast, Known Unknowns are information that models correctly recognize as unknowns, which is essential for developing trustworthy AI. A model’s self-awareness capability is directly proportional to its grasp of information within the Known Knowns and Known Unknowns quadrants. It is crucial for models to identify their limitations in processing information to avoid providing incorrect responses, a consideration existing benchmarks have often overlooked. Thus, in the following sections, we detail our approach to constructing data that assesses the self-awareness of MLLMs according to the proposed quadrant.
3.2 MM-SAP Benchmark
To evaluate the self-awareness of MLLMs, we proposed the MM-SAP benchmark, consisting of three VQA datasets that respectively correspond to the previously mentioned categories of information. We provides a comprehensive overview in Figure 3, illustrating the sub-datasets of MM-SAP along with their respective proportions. Additionally, Figure 4 displays examples from each sub-datasets. More detailed statistics of the dataset can be found in Appendix A.1. In this section, we introduce the construction of the three individual sub-datasets in detail.
Figure 4: Examples for each sub-dataset. In MM-SAP, all samples include a refusal option. In BeyondVisQA, the model can only choose the refusal option. In KnowVisQA, the model has the option to select either the correct answer or to correctly refuse to answer. In BasicVisQA, the model is restricted to choosing the correct option only.
BasicVisQA
Basic Visual Information QA (BasicVisQA) is specifically designed to evaluate the model’s self-awareness capability, particularly in ‘known knowns’. This dataset includes questions that cover eight types of basic visual information, as illustrated in Figure 3, such as coarse-grain object recognition and color recognition. As previously discussed, these information categories are all considered ‘knowns’ to MLLMs. To construct BasicVisQA, we sampled questions from the VQAv2  Goyal et al. (2017) validation set that pertained to basic visual information. To increase the dataset’s complexity, we manually crafted additional 150 questions using images sourced from COCO  Lin et al. (2014) and Visual Genome  Krishna et al. (2017). Moreover, for each question, we generated three incorrect yet plausible options alongside the correct one. We also introduced a refusal option for each question, as depicted in Figure 4, allowing the model to opt out of answering. Consequently, BasicVisQA comprises 400 questions accompanied by 397 images, with each question offering five distinct choices.
KnowVisQA
Knowledge-intensive Visual Information QA (KnowVisQA) consists of perceptual questions that require visual knowledge for answering. We focus on six distinct domains as illustrated in Figure 3: animals and plants, brands and products, art, landmarks, food, and organizations. Images for these domains were collected from various online sources, followed by the meticulous formulation of 350 questions, each accompanied by five options, as seen in Figure 4. Unlike previous knowledge-based VQA datasets such as OKVQA Marino et al. (2019) or A-OKVQA Schwenk et al. (2022), KnowVisQA focus on visual knowledge and incorporates a refusal option for evaluation.
BeyondVisQA
We have developed a novel VQA dataset named Beyond Visual Information QA (BeyondVisQA), This dataset is specifically designed to assess the ‘known unknowns’ self-awareness capability of a MLLM. It includes questions that require information beyond what the input images provide. We have divided these questions into six distinct categories, as shown in Figure 3. The details of the categories are provided in Appendix A.2.We meticulously crafted 400 unanswerable questions based on a sample of 308 images from the COCO and Visual Genome datasets. Additionally, for each question, we generated four plausible yet misleading options along with one refusal option. This dataset serves as a crucial component in assessing the self-awareness capabilities of various MLLMs regarding ‘known unknowns’. It helps measure their ability to identify information beyond what is visible in images.
Model BasicVisQA KnowVisQA BeyondVisQA Total
$score_{kk}$ $score_{kk}$ $score_{ku}$ $score_{ku}$ $score_{kk}$ $score_{ku}$ $score_{sa}$
LLaVA-7b 60.75 46.06 1.37 25.70 35.15 9.36 44.50
LLaVA-13b 66.35 48.86 1.49 30.85 37.95 11.18 49.13
instructblip-vicuna-7b 70.10 46.17 4.11 38.05 38.43 14.49 52.92
InfMLLM-7b 70.10 46.17 4.11 38.05 38.43 14.49 52.92
InternLM-XComposer2-VL-7b 73.05 53.49 0.74 37.55 41.69 13.29 54.97
Yi-VL-6B 60.65 52.74 5.49 25.25 37.15 10.45 47.60
ShareGPT4V-7b 65.80 48.51 1.83 36.80 37.65 13.36 51.01
ShareGPT4V-13b 66.30 51.89 0.80 25.75 38.85 9.20 48.05
CogVLM-17b 65.20 61.66 0.69 29.85 41.44 10.59 52.03
Qwen-VL-Chat-7b 62.15 63.31 1.43 18.90 40.89 7.01 47.90
Qwen-VL-Plus* 70.50 71.71 2.86 63.50 46.35 24.18 70.53
Qwen-VL-Max* 75.00 78.00 3.77 70.25 49.83 25.58 75.41
Gemini 1.0 Pro Vision* 62.75 70.85 1.71 52.25 43.49 18.69 62.18
GPT-4V* 63.20 63.60 12.06 77.25 41.34 30.54 71.88
Table 1: Overall results of various MLLMs on MM-SAP. We present only the value of $score_{kk}$ for BasicVisQA, as the questions within it are all known for MLLMs. Similarly, we only display the value of $score_{ku}$ for BeyondVisQA. Bold values indicate the highest mean score in each column. Closed-source MLLMs are marked with ’*’.
4 Experiments
4.1 Evaluation Strategy
Self-awareness encompasses the abilities to recognize ‘knowns’ and ‘unknowns’. Accordingly, we introduce three metrics to measure a model’s self-awareness in the MM-SAP benchmark.
  • $score_{kk}$: It represents the proportion of the question answer correctly by the model.
  • $score_{ku}$: It represents the proportion of questions that the model correctly rejects.
  • $score_{sa}$: It is the sum of $score_{kk}$ and $score_{ku}$, representing the self-awareness of a model.
Figure 5: Scores distribution of MLLMs. The x-axis and y-axis represent the $score_{kk}$ and $score_{ku}$ respectively. The dashed lines in the figure represent the isoline of the $score_{sa}$.
Before describing the calculation of the above metrics, we first define some indicators to avoid confusion. For each question $q_{i}$ in the test set $\bm{q}$, we denote the indexes of the correct option and the refusal option as $c_{i}$ and $r_{i}$, respectively. Note that $c_{i}$ for $q_{i}\in\bm{q}_{\rm beyond}$ does not exist. Therefore, $score_{kk}$ and $score_{ku}$ can be defined as:
${\rm score}_{kk}$ $=\frac{100\cdot\sum^{|\bm{q}|}_{i=1}\mathbb{I}(p_{i}=c_{i})\cdot\mathbb{I}(q_{i}\text{ is known})}{|\bm{q}|}$ (1)
$=\frac{100\cdot\sum^{|\bm{q}|}_{i=1}\mathbb{I}(p_{i}=c_{i})}{|\bm{q}|}$
${\rm score}_{ku}=\frac{100\cdot\sum^{|\bm{q}|}_{i=1}\mathbb{I}(p_{i}=r_{i})\cdot\mathbb{I}(q_{i}\text{ is unknown})}{|\bm{q}|}$ (2)
where $p_{i}$ represents the prediction of the evaluated MLLM for $q_{i}$. We omit the term $\mathbb{I}(q_{i}\text{ is known})$ in Equation 1 because the questions that model can correctly answer are all considered ‘knowns’.
For $q_{i}$ in BasicVisQA and BeyondVisQA, determining the value of $\mathbb{I}(q_{i}\text{ is unknown})$ is straightforward because they are respectively ‘knowns’ and ‘unknowns’ for models. For $q_{i}\in\bm{q}_{know}$, the condition $p_{i}=r_{i}$ does not necessarily imply that $q_{i}$ is unknown, as models might refuse to answer questions they actually know. To address this issue, we remove the refusal option and compel the model to choose an answer. If the model selects the correct one, it indicates that the model actually knows the answer. Consequently, $\mathbb{I}(q_{i}\text{ is unknown})$ can be defined as follows:
$\mathbb{I}$ $(q_{i}\text{ is unknown})=$ (3)
$\begin{cases}0&\text{if }q_{i}\in\bm{q}_{basic},\\ \mathbb{I}(p_{i}^{\prime}\neq c_{i}\mid p_{i}=r_{i})&\text{if }q_{i}\in\bm{q}_{know},\\ 1&\text{if }q_{i}\in\bm{q}_{beyond}\end{cases}$
where $p_{i}^{\prime}$ is the model’s prediction without the refusal option. The self-awareness score($score_{sa}$) is then calculated as:
${\rm score}_{sa}$ $={\rm score}_{kk}+{\rm score}_{ku}$ (4)
Model BasicVisQA KnowVisQA BeyondVisQA
Answer Rate Answer Acc Answer Rate Answer Acc Answer Rate
LLaVA-7b 98.70% 61.55% 98.46% 46.78% 74.30%
LLaVA-13b 99.10% 66.95% 97.60% 50.05% 69.15%
InfMLLM-7b 98.35% 71.28% 92.86% 49.72% 61.95%
InternLM-XComposer2-VL-7b 99.45% 73.45% 98.86% 54.10% 62.45%
Yi-VL-6B 98.10% 61.83% 91.89% 57.41% 74.75%
ShareGPT4V-7b 97.60% 67.42% 97.54% 49.74% 63.20%
ShareGPT4V-13b 99.10% 66.10% 98.57% 52.63% 74.25%
CogVLM-17b 98.85% 65.96% 98.97% 62.30% 70.15%
Qwen-VL-Chat-7b 97.40% 63.81% 99.71% 63.50% 81.10%
Qwen-VL-Plus* 98.25% 71.76% 96.86% 74.04% 36.50%
Qwen-VL-Max* 97.95% 76.57% 96.91% 80.48% 29.75%
Gemini 1.0 Pro Vision* 99.00% 63.38% 97.13% 72.86% 47.75%
GPT-4V* 94.45% 66.90% 83.83% 75.87% 22.75%
Table 2: Results of Answer Rate and Answer Accuracy of MLLMs on MM-SAP. Except for the Answer Rate in BeyondVisQA, where a lower rate is better, higher values indicate better performance for all other metrics. Bold numbers highlight the best mean value in each column. Models marked with ’*’ are closed-source.
4.2 Inference Settings
For all the MLLMs tested in this study, we set the decoding temperature to $t=0$ and the decoding beam size to $b=1$. To reduce the uncertainty of the scores, each model is requested to predict the answer five times, with the order of the options randomly shuffled. We then calculate the mean of all scores as the result. More evaluation details are provided in Appendix B.
4.3 Main Results
A total of thirteen popular MLLMs were evaluated on our MM-SAP benchmark, including LLaVA-7B, LLaVA-13B Liu et al. (2023b, c), ShareGPT4V-7B, ShareGPT4V-13B Chen et al. (2023), CogVLM-17B Wang et al. (2023b), Yi-VL-6B Yi (2023), Qwen-VL-Chat, Qwen-VL-Plus, Qwen-VL-Max Bai et al. (2023), InfMLLM-7B Zhou et al. (2023), InternLM-XComposer2-VL-7B Dong et al. (2024), Gemini 1.0 Pro Vision Team et al. (2023), and GPT-4V OpenAI (2023b). The self-awareness scores $score_{sa}$ of these MLLMs are presented in Table 1.
As shown in Table 1 and Figure 5, there is a significant difference in the $score_{sa}$ between closed-source and open-source MLLMs. Qwen-VL-Max achieves the highest $score_{sa}$, with the other two closed-source models also scoring closely, significantly outperforming open-source models. In terms of ‘known knowns’, Qwen-VL-Plus and Qwen-VL-Max achieve high $score_{kk}$ on both BasicVisQA and KnowVisQA, while GPT-4V does not show obvious advantage compared to open-source models. When it comes to $score_{ku}$, however, GPT-4V demonstrates particularly notable performance. In BeyondVisQA, the proportion of correctly refused questions by open-source models does not exceed 40%, while closed-source models reach up to 70%. The ability to recognize unknowns—information not provided in the images—among Qwen-VL-Plus, Qwen-VL-Max, and GPT-4V is relatively similar. However, only GPT-4V clearly demonstrates the ability to refuse to answer questions beyond its intrinsic visual knowledge. This is evident in KnowVisQA, where GPT-4V’s $score_{ku}$ of 12.06% significantly surpasses those of the other models, indicating GPT-4V’s superior awareness of its visual knowledge boundaries. Despite a lower $score_{sa}$ compared to Qwen-VL-Max, GPT-4V’s ability to identify ‘unknowns’ is distinctly superior.
Model BasicVisQA KnowVisQA BeyondVisQA Total
$score_{kk}$ $score_{kk}$ $score_{ku}$ $score_{ku}$ $score_{kk}$ $score_{ku}$ $score_{sa}$
InfMLLM-7b 70.10 46.17 4.11 38.05 38.43 14.49 52.92
InfMLLM-7b + prompt 64.90 42.06 10.63 56.35 35.37 22.83 58.21
ShareGPT4V-7b 65.80 48.51 1.83 36.80 37.65 13.36 51.01
ShareGPT4V-7b+prompt 64.70 48.06 3.03 41.30 37.13 15.29 52.42
GPT-4V* 63.20 63.60 12.06 77.25 41.34 30.54 71.88
GPT-4V*+prompt 58.85 59.20 16.86 87.00 38.49 35.39 73.88
Table 3: Results of the prompting strategy. Bold values indicate the highest mean score in each column. Closed-source MLLMs are marked with ’*’
4.4 Refusal Behavior of MLLMs
To provide a more comprehensive analysis, we define the following two indicators to study the models’ refusal behavior.
${\rm Answer\ Acc}=\frac{\sum^{|\bm{q}|}_{i=1}\mathbb{I}(p_{i}=\bm{c}_{i})}{\sum^{|\bm{q}|}_{i=1}\mathbb{I}(p_{i}\neq r_{i})}$ (5)
${\rm Answer\ Rate}=\frac{\sum^{|\bm{q}|}_{i=1}\mathbb{I}(p_{i}\neq r_{i})}{|\bm{q}|}$ (6)
where the Answer Accuracy is the proportion of the correct predictions among the questions that answered, and the Answer Rate is the proportion of all questions that the model attempts to answer.
Table 2 presents the results for the Answer Rate and Answer Accuracy of MLLMs. The results reveal that the Answer Rates for most open-source models on BasicVisQA and KnowVisQA are nearly 100%. GPT-4V exhibits the lowest Answer Rate, indicating its superior ability to recognize what it does not know. Additionally, it is noted that GPT-4V incorrectly rejects some questions in BasicVisQA, suggesting that its tendency towards refusal somewhat impacts its ability to process known information. For KnowVisQA, GPT-4V exhibits the lowest Answer Rate, highlighting its capability to decline answering some unknown questions and avoide generate incorrect responses.
To delve deeper into the refusal behavior on KnowVisQA, we selected four models with relatively low Answer Rates for further analysis. We define the following two indicators:
${\rm Refusal\ Num}={\sum^{|\bm{q_{know}}|}_{i=1}\mathbb{I}(p_{i}=\bm{r}_{i})}$ (7)
${\rm Unknown\ Knowns\ Rate}=$ (8)
$\quad\quad\quad\frac{\sum^{|\bm{q_{know}}|}_{i=1}\mathbb{I}(p_{i}=\bm{r}_{i})\cdot\mathbb{I}(p_{i}^{\prime}=\bm{c}_{i})}{|\bm{q_{know}}|}$
Table 4 shows that the Unknown Knowns Rate for InfMLLM-7b is 42.47%, indicating that nearly half of the questions it refused were actually known to it. While Qwen-VL-Max exhibits the lowest Unknown Knowns Rate, its Refusal Number is comparatively low. GPT-4V has the highest Refusal Number and a relatively low Unknown Knowns Rate, suggesting its capability to refuse some unknown questions. However, considering the Answer Accuracy detailed in Table 2, we observe that current models struggle to accurately identify unknown visual knowledge, indicating significant room for improvement.
Model Refusal Num Unknown Knowns Rate
InfMLLM-7b 25.0 42.47%
Yi-VL-6b 28.4 32.10%
Qwen-VL-Max* 10.8 14.27%
GPT-4V* 56.6 26.19%
Table 4: Results of the Refusal Num and the Unknown Knowns Rate of MLLMs. Closed-source MLLMs are marked with ’*’.For each MLLM, we conducted five experiments and report the mean result, which explains why the Refusal Num is not an integer.
4.5 Recognizing Unknows through Prompting
Given the capability of many MLLMs to follow instructions, we attempted to directly instruct an MLLM to choose the refusal option when confronted with unknown questions by appending a prompt to the text input. This prompt, termed the ‘refusal prompt’, is as follows: “Answer with the option’s letter from the given choices directly. If you don’t know the answer, please reply with ‘Sorry, I can’t help with it’.”. Experiments were conducted on three MLLMs with relatively high $score_{ku}$ , to evaluate the effectiveness of this prompting strategy.
Table 3 demonstrates the comparative results before and after using the refusal prompt. The introduction of the refusal prompt notably improves the $score_{ku}$, yet the scores on KnowVisQA remain considerably low. Additionally, the refusal prompt negatively affects $score_{kk}$. Therefore, the application of simple prompting strategy results in limited improvement in the model’s $score_{sa}$, indicating the necessity for further research to effectively enhance the self-awareness capabilities of MLLMs.
5 Conclusion
In this paper, we introduce MM-SAP, a novel benchmark designed to evaluate self-awareness in perception for MLLMs. By innovatively integrating image information with knowledge quadrants, we have developed a modified quadrant specifically tailored for MLLMs. Building on this, we present the MM-SAP benchmark, which comprises three distinct sub-datasets. We conducted evaluations of various MLLMs using this benchmark and analyzed their results to gain insights into the self-awareness capabilities of these models. We believe that the MM-SAP benchmark offers a nuanced and detailed perspective on the self-awareness of MLLMs, contributing significantly to the development of more trustworthy and reliable AI systems.
6 Limitations
In our study, we specifically assess self-awareness in perception, omitting the more intricate cognitive tasks. While these aspects are significant, they introduce complexity into data collection and analysis. Furthermore, the proposed MM-SAP benchmark comprises only multiple-choice problems. However, the actual application scenarios for MLLMs typically involve open-ended questions and interactions. Providing models with options could potentially give them hints and simplify the task’s complexity, thereby resulting in an overestimation of the models’ self-awareness compared to their performance in real-world applications.
Acknowledgments
This work is supported by National Key R&D Program of China (No. 2022ZD0162101), STCSM (No. 21511101100, No. 22DZ2229005), and State Key Laboratory of UHD Video and Audio Production and Presentation.
Appendix A MM-SAP benchmark
We provide more details on the MM-SAP benchmark.
A.1 Statistic of Dataset
The average number of words in queries and options is presented in Table 5. Additionally, Table 6 shows the distribution of correct and refusal options.
average number of words
queries 9.8
options 2.8
Table 5: The average number of words in querie and options in MM-SAP.
A B C D E
correct options 20.3% 19.1% 18.8% 20.6% 21.2%
refusal options 21.6% 20.3% 18.7% 19.9% 19.5%
Table 6: The choice distribution of correct options and refusal options in MM-SAP.
A.2 The Categories of Questions in BeyondVisQA
BeyondVisQA encompasses six distinct categories of questions as follows:
  • Nonexistent Objects: Questions about elements not present in the image, requiring inference beyond the visual information provided.
  • Background Information: Questions that seek background details about objects not depicted in the image.
  • Temporal Unpredictability: Questions about events or conditions that occurred before or after the moment captured in the image.
  • Missing Visual Information: Questions about details that are visually unclear, hidden, or blurred in the image.
  • Other Modalities Information : Questions that require information from non-visual modalities, such as sound or smell, which images cannot convey.
  • Intractable Quantity: Questions that involve quantifying elements that cannot be accurately determined from the image’s visual information alone.
All these questions are considered unknowns because they require information beyond the image provided to be answered.
Appendix B Evaluation Detail
The prompts used for model evaluation are shown in Figure 6. To evaluate the responses, we employed two methods: calculating the perplexity (PPL) and directly matching the characters of the options. We applied these methods to LLaVA and ShareGPT4V as shown in table 7 and found that the results were nearly identical. Given that closed-source models like GPT-4V and Qwen-VL-Max cannot be evaluated using PPL calculations, we ultimately decided to evaluate answer correctness by directly matching the characters of the options for all models.
Figure 6: Prompts for model evaluation
Model BasicVisQA KnowVisQA BeyondVisQA Total
$score_{kk}$ $score_{kk}$ $score_{ku}$ $score_{ku}$ $score_{kk}$ $score_{ku}$ $score_{sa}$
LLaVA-7b(direct matching) 60.75 46.06 1.37 25.70 35.15 9.36 44.50
LLaVA-7b(PPL) 60.75 46.36 1.37 25.80 35.23 9.40 44.63
ShareGPT4V-7b(direct matching) 65.80 48.51 1.83 36.80 37.65 13.36 51.01
ShareGPT4V-7b(PPL) 65.75 48.71 1.83 37.05 37.68 13.45 51.14
Table 7: The results of LLaVA-7b and ShareGPT4V-7b with different evaluating method.
Appendix C Additional Examples in MM-SAP
In this section, we provide supplementary examples from our MM-SAP benchmark as shown in Figure 7, Figure 8, and Figure 9.
Figure 7: Supplementary Examples in BasicVisQA.
Figure 8: Supplementary Examples in KnowVisQA
Figure 9: Supplementary Examples in BeyondVisQA